OpenStudy (anonymous):

Can someone explain why |x| = sqrtx^2

OpenStudy (anonymous):

\[\sqrt{9} = \pm3\]"sqrt of positive is also always positive" ???

OpenStudy (experimentx):

sorry about that

OpenStudy (anonymous):

so, anyone???

OpenStudy (experimentx):

then the above relation is not true

OpenStudy (experimentx):

i mean not absolutely

OpenStudy (anonymous):

thats what i think, but someone just said it is, turinTest, i think. And he was preaty sure about it....

OpenStudy (experimentx):

can you show me where??

OpenStudy (experimentx):

√9 !=-3 but √9=+3

OpenStudy (anonymous):

ya, i just so his explanation, but loks so forced to me. Maybe just my problem

OpenStudy (experimentx):

yeah that's right, because he is taking the positive root of x, which is equal to the positive value of x

OpenStudy (experimentx):

but x can one value between -ve and +ve, but absolute value of x will always be equal to the positive root of x^2

OpenStudy (anonymous):

\[|x|=-x (x<0)\]\[|x| = x (x>0)\] so, to puting them bouth together he using |x| for root values

OpenStudy (anonymous):

root is multivalued function...

OpenStudy (experimentx):

to be honest I made silly statement with out a thought. 'root of positive is always positive' ... but i should rather have been 'positive root of ..'

OpenStudy (anonymous):

right

OpenStudy (anonymous):

When you square a number, you lose the sign. Once it is gone, it doesn't come back. So taking the root of the square is just like taking the absolute value.

OpenStudy (experimentx):

he is taking THE POSITIVE ROOT

OpenStudy (experimentx):

heheh ... positive root will always be positive and negative root will always be negative. in above case you are also taking the positive root

OpenStudy (anonymous):

hmm.. I agree with sign los, but when you take the root you will get +/- anyway

OpenStudy (experimentx):

and also -|x| = - sqrtx^2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so i guess:\[|x| = \sqrt{x ^{2}}\] is just the way to write: \[\pm x = \sqrt{x ^{2}}\]

OpenStudy (experimentx):

no ... the second is not right, because on the right hand side ... you have already chosen the positive root

OpenStudy (anonymous):

i don't think i chouse anything. x can be negative, can be positive. Same for the sqrt

OpenStudy (experimentx):

the EXPRESSION on right hand side must be positive

OpenStudy (anonymous):

x2 is >0 no mater what x is

OpenStudy (experimentx):

yup, that is if x has real value

OpenStudy (anonymous):

sorry, my browser won't let me use the equation applet, but: (the square root of x) squared is always going to be positive

OpenStudy (anonymous):

and why? plz explain

OpenStudy (anonymous):

any number squared will be positive, no?

OpenStudy (experimentx):

because if x has a real value, then the value of x is either going to be positive or negative .. and the positive multiplication with positive is always positive and -x- is also always postive

OpenStudy (anonymous):

if real yes. But continue

OpenStudy (experimentx):

But??

OpenStudy (anonymous):

I have no imagination. Nor do I understand sarcasm. Sorry.

OpenStudy (anonymous):

my point is: you "aplye" srt function to a square of a nummber. Root function gives you two possible values +/-. You don't know what x (+,-) was used to get the square.

OpenStudy (experimentx):

the root function doesn't give you two value ... the root function gives you one value ... it's only becuse you have two roots, one negative the other positve

OpenStudy (experimentx):

and also be definition of function, function gives you one value ..

OpenStudy (anonymous):

there are multivalued functions, and square root is one of them...

OpenStudy (anonymous):

arsen another

OpenStudy (anonymous):

arcsen*

OpenStudy (experimentx):

let me repeat this, a function CANNOT give you two values,

OpenStudy (anonymous):

it can and it gives

OpenStudy (experimentx):

and arcsin(x) is not a function, if it's range is not defined with in [-pi/2, pi/2]

OpenStudy (experimentx):

In mathematics, a function[1] is a correspondence from a set of inputs to a set of outputs that associates each input with exactly one output --- from wikipedia

OpenStudy (anonymous):

a function is a set of ordered pairs of one, or two sets. No more restrictions are aplyed

OpenStudy (accessdenied):

a function maps each value in its domain to one not necessarily unique value of the range you will not have a function by definition with multiple y-values for one x-value

OpenStudy (experimentx):

no .. not exactly, a function must give exactly one output for one or many input // as accessdenied said, unique is not necessary condition ... but it must give one output ... not many outputs

OpenStudy (anonymous):

agree about that each value of the domain is mapped. But can be mapped to more than one value in codomain

OpenStudy (anonymous):

sqrt of 1 has 2 values. Third root: 3, and so on

OpenStudy (anonymous):

nth root has n values

OpenStudy (experimentx):

no .. sqrt has one and only one value X^2 - 1 = 0, x=+-1 are roots of equation on/// not the value of srt 1

OpenStudy (anonymous):

sqrt has two values. The positive one is cold principal value, but there is also a negative one.

OpenStudy (experimentx):

can you show me in graph???

OpenStudy (experimentx):

f(x) = sqrt(x)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Multivalued_function look there in examples for multivalued functions. For graphing you need to deside which value will you use, normaly principal is taken

OpenStudy (anonymous):

sqrt graph maps positive R to it self, but that principal value of the root

OpenStudy (experimentx):

multivalued function and function are different things ... they have different criteria and different things ,, the graph of sqrt x is given here http://www.wolframalpha.com/input/?i=%28x%29%5E1%2F2

OpenStudy (anonymous):

i know the graph, thats not the point of my question

OpenStudy (experimentx):

if sqrt x is a multivariable function then it must be something like this